Last month when Holiday Creations, Inc., sold 41,000 units, total sales were $306,000, total variable expenses were $253,980, and fixed expenses were $38,600.


Required:
1.

What is the company’s contribution margin (CM) ratio?

     

2.

Estimate the change in the company’s net operating income if it were to increase its total sales by $1,900.
Answer #1

Contribution margin=Sales-Variable cost  

=(306,000-253980)=$52020

a.Contribution margin ratio=Contribution margin/Sales

=(52020/306000)=17%

b.Change in net operating income=Increase in net operating income

=Increase in Sales*Contribution margin ratio

=1900*17%

=$323